Decide if these equations represent parallel lines, perpendicular lines, or neither
4x+8y=12
y=2x+10


y=3x+2
5x+2y=20


y=3x+2
y=3/1x+15

Answer by Alan3354(30993)

Decide if these equations represent parallel lines, perpendicular lines, or neither
===============
Solve for y.
Then the slope, m, = the coefficient of x
=================================================
4x+8y=12
y=2x+10
--------------
Solve for y
4x+8y=12x
8y = 8x
y = x
Slope = 1
-----
y=2x+10
Slope = 2
If the slopes are equal, they're parallel.
If the slopes are negative inverses, they're perpendicular.
These 2 are:
Neither
-----------------
Do the others the same way.
y=3x+2
5x+2y=20


y=3x+2
y=3/1x+15